# Uncovering Quantum Many-body Scars with Quantum Machine Learning

### v1

@@ V.2 Approximate quasiparticle states
Figure 8: (color online) The mean and variance of the total $S^{z}$ for each eigenstate within a certain $n_{\rm dw}$ sector: (a)&(c) $n_{\rm dw}=2$; (b)&(d) $n_{\rm dw}=3$. The red crosses indicate the eigenstates identified by the QCNN. The green solid curves represent analytical results from the ferromagnetic magnon bound state approximation, while the purple dashed curves correspond to those obtained from the anti-ferromagnetic magnon bound state approximation. Other parameters are $\lambda=J=2\Delta$.
Figure 9: (color online) Illustration of two types of quasiparticles moving via intermediate processes. On the left side of the dashed line, blue spins form a ferromagnetic background for the propagation of the bound state. On the right side of the dashed line, blue spins are trapped in an antiferromagnetic configuration. (a) The motion of a ferromagnetic bound state (red spins). (b) The motion of an anti-ferromagnetic bound state (red spins).
V.2.1 Ferromagnetic magnon bound state
@@
The analytical solution of the ground state energy is $E_{k}=-\Delta\sqrt{10+3u^{2}+\sqrt{64+48u^{2}+5u^{4}}}/\sqrt{2}$, where $u=\left|\lambda+\lambda e^{ik}\right|/\Delta$. Under OBC, the system approximately forms standing waves as a superposition of states with momenta ±$\pm k$. By varying $k$, we calculate $E_{k}$ and the corresponding $S_{z}$. The relationship between these quantities is illustrated by the green solid curves in Figs. 7 and 8. Most of the marked states closely align with these curves, suggesting that they can indeed be interpreted as spin-wave modes. Furthermore, as $\Delta$ increases, the agreement between the marked states and the analytical approximation improves, as shown in Fig. 8.
Figure 10: (color online) Weights of each eigenstate in the spin-wave mode of (a)(b) ferromagnetic magnon bound state, (c)(d) anti-ferromagnetic magnon bound state. The special states marked by QCNN have anomalously large weight on the subspace of a particular type of quasiparticle, compared to typical eigenstates. Other parameters are $\lambda=J=10\Delta$, (a)(c)$n_{\rm dw}=2$, (b)(d)$n_{\rm dw}=3$.
V.2.2 Anti-ferromagnetic magnon bound state
@@
The energy of the ground state and the first excited state are $E_{k}=-\sqrt{\Delta^{2}+2\lambda^{2}(1\pm\cos(k/2))}$. Since the energy of the first excited state is closer to the exact scar state, we present it with purple dashed curves in Figs. 7 and 8. The deviation of the antiferromagnetic magnon bound states from the exact eigenstates is greater than that of the ferromagnetic magnon bound states, leading to a smaller component contribution, as shown in Fig. 10.
@@
VI Experimental demonstration on quantum device
@@
Figure 11: (color online) Success rate of the QCNN during training on a quantum device. The experiment was conducted in 12 groups, each spaced hours apart to assess drift error. Each group involved $10^{4}$ measurement repeats to determine the success probability. Error bars indicated by the shaded region represent the standard error across different groups.
@@
Figure 12: (color online) (a)&(b): Two different methods for boosting errors in the circuit: (a) adding single-qubit Pauli errors; (b) replacing $U$ with $U\left(U^{\dagger}U\right)^{r}$. (c) Error mitigation according to input fidelity. Linear regression suggests $P_{1}\approx 67\pm 1\%$ in the limit of a perfect input state. The red and blue dots correspond to Monte Carlo sampling of single-qubit Pauli errors (averaged over $10^{3}$ noise realizations) and repeating $U^{\dagger}U$, respectively. (d) Error mitigation according to error rate in the second approach. Linear regression using data prior to saturation suggests $P_{1}\approx 63\pm 1\%$ in the limit of a perfect input state.
VII Generalizations to other models
@@
Figure 13: (color online) (a) Overlap between the eigenstates and the $|Z_{2}\rangle$ state, with red crosses marking the states identified by the QCNN. (b) Density of QCNN-identified states based in (a). (c) Density of energy differences between the QCNN-identified states in (a). In both (b) and (c), only states with non-zero overlap with the $|Z_{2}\rangle$ state are counted, with smoothing applied using Gaussian broadening of each points set to $20.1\Omega/$ 2.
Figure 14: (color online) (a) Overlap between the eigenstates and the $Z_{1001}$ state. The red crosses are states marked by QCNN. (b) Density of QCNN-identified states based in (a). (c) Density of energy differences between the QCNN-identified states in (a). In both (b) and (c), only states with non-zero overlap with the $|Z_{1001}\rangle$ state are counted, with smoothing applied using Gaussian broadening of each points set to $0.25J_{\rm e}$.
@@
In this section, we extend our QCNN-based approach to two additional models that host QMBSs. We begin with the PXP model, utilizing training data that include both analytically solvable and numerically identified scar states. Specifically, we incorporate the four exact scar states reported in Ref. [30] and consider the states with the largest overlap with the $|Z_{2}\rangle$ across various energy windows as scar states [29]. Notably, the QCNN identifies many additional states as potentially non-thermal states, as shown in Fig. 13(a). The dynamics of these states are governed by their energy spectrum. Although some of these states have smaller overlaps with $|Z_{2}\rangle$, they share similar energies with the $Z_{2}$ tower states. These states form distinct energy towers, as depicted in Fig. 13(b). Their energies are approximately equally spaced, as shown in Fig. 13(c), indicating that their linear superposition can lead to stable oscillations.
@@
VIII Discussions and conclusion
Recent advances in the use of QML to investigate scar states have garnered considerable interest, resulting in a variety of related research. Here, we discuss differences between their studies and our approach. For instance, Ref. [26] introduces a general classifier using quantum variational autoencoders, where each eigenstate is assigned its own learning circuit. This approach can be computationally expensive, with costs scaling exponentially as Hilbert spaces grow in size. In contrast, our method employs a single circuit to learn the common properties of known scar states, enabling the discovery of other similar states. This makes our scheme more computationally feasible. Meanwhile, Ref. [44] employs classical machine learning to detect single-body scar in billiard systems, given the classical description of eigenstates, but this methodology does not extend to many-body scars. Ref. [27] uses classical machine learning and dimensionality reduction techniques to minimize the distance among scar states, but this can lead to loss of quantum information.
@@ VIII Discussions and conclusion
Acknowledgments
J.J.F., B.Z. and Q.Z. acknowledge support from University of Southern California. Z.-C. Y. acknowledges support from Peking University. This research was conducted using IBM Quantum Systems provided through USC’s IBM Quantum Innovation Center.
@@ Appendix A Details on the QCNN
$U|A_{i}\rangle$ $=$ $|a_{i}\rangle\otimes|0\rangle~{},$ (11)
$U|B_{j}\rangle$ $=$ $|b_{j}\rangle\otimes|1\rangle~{},$ (12)
@@
$U|A_{i}\rangle\otimes|0\rangle$ $=$ $|\alpha_{i}\rangle\otimes|0\rangle~{},$ (13)
$U|B_{j}\rangle\otimes|0\rangle$ $=$ $|\beta_{j}\rangle\otimes|1\rangle~{},$ (14)
@@ Appendix B Details of experiments
Several methods exist for generating exact scar states [45, 28]. For producing the first exact scar state, we employ a more convenient approach. The procedure for generating the first scar state is outlined in Fig. 17. All the qubits are initialized to $|0\rangle$. The last qubit $q_{n-1}$ is flipped to $|1\rangle$ using a not gate (X). An equal superposition is created on the ancilla qubit with a Hadamard gate (H). This is followed by a controlled-SWAP gate, which generates a superposition of spins at different positions $\left(|0_{\rm a}\rangle\otimes|100\cdots\rangle+|1_{\rm a}\rangle\otimes|\cdots 010\cdots\rangle\right)/\sqrt{2}$. Next, a CNOT gate is applied to restore the ancilla qubit, resulting in $|0_{\rm a}\rangle\otimes\left(|100\cdots\rangle+|\cdots 010\cdots\rangle\right)/\sqrt{2}$. Uniform distribution along the chain requires $O(n)$ equal divisions. Finally, a Z gate is applied to adjust the phase.
@@
${\rm Tr}\left\{|\mathscr{S}_{1}\rangle\langle\mathscr{S}_{1}|\rho_{\varepsilon}\right\}$ $=$ ${\rm Tr}\large\{U|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|U^{\dagger}U_{\varepsilon}|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|U_{\varepsilon}^{\dagger}\large\}~{},$
@@
$X_{n-1}U^{\dagger}U_{\varepsilon}X_{n-1}\approx U^{\dagger}U_{\varepsilon}~{}.$ (16)
@@
$\approx$ ${\rm Tr}\large\{|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|H_{\rm s}U_{\varepsilon}X_{n-1}|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|X_{n-1}U_{\varepsilon}^{\dagger}H_{\rm s}\large\}~{}.$

### v2

@@ II.3.2 Approximate quasiparticle states
Figure 7: The mean and variance of the total $S^{z}$ for each eigenstate within a certain $n_{\rm dw}$ sector: (a)&(c) $n_{\rm dw}=2$; (b)&(d) $n_{\rm dw}=3$. The red crosses indicate the eigenstates identified by the QCNN. The green solid curves represent analytical results from the ferromagnetic magnon bound state approximation, while the purple dashed curves correspond to those obtained from the anti-ferromagnetic magnon bound state approximation. Other parameters are $\lambda=J=2\Delta$.
Figure 8: Illustration of two types of quasiparticles moving via intermediate processes. On the left side of the dashed line, blue spins form a ferromagnetic background for the propagation of the bound state. On the right side of the dashed line, blue spins are trapped in an antiferromagnetic configuration. (a) The motion of a ferromagnetic bound state (red spins). (b) The motion of an anti-ferromagnetic bound state (red spins).
II.3.3 Ferromagnetic magnon bound state
@@
The analytical solution of the ground state energy is $E_{k}=-\Delta\sqrt{10+3u^{2}+\sqrt{64+48u^{2}+5u^{4}}}/\sqrt{2}$, where $u=\left|\lambda+\lambda e^{ik}\right|/\Delta$. Under OBC, the system approximately forms standing waves as a superposition of states with momenta ±$\pm k$. By varying $k$, we calculate $E_{k}$ and the corresponding $S_{z}$. The relationship between these quantities is illustrated by the green solid curves in Figs. 6 and 7. Most of the marked states closely align with these curves, suggesting that they can indeed be interpreted as spin-wave modes. Furthermore, as $\Delta$ increases, the agreement between the marked states and the analytical approximation improves, as shown in Fig. 7.
Figure 9: The spin-wave component in each eigenstate. The weights of the ferromagnetic magnon bound state are shown in (a) and (b), while the weights of the anti-ferromagnetic magnon bound state are shown in (c) and (d). The special states marked by QCNN have anomalously large weight on the subspace of a particular type of quasiparticle, compared to typical eigenstates. Other parameters are $\lambda=J=10\Delta$, (a)(c)$n_{\rm dw}=2$, (b)(d)$n_{\rm dw}=3$.
II.3.4 Anti-ferromagnetic magnon bound state
@@
The energy of the ground state and the first excited state are $E_{k}=-\sqrt{\Delta^{2}+2\lambda^{2}(1\pm\cos(k/2))}$. Since the energy of the first excited state is closer to the exact scar state, we present it with purple dashed curves in Figs. 6 and 7. The deviation of the antiferromagnetic magnon bound states from the exact eigenstates is greater than that of the ferromagnetic magnon bound states, leading to a smaller component contribution, as shown in Fig. 9.
@@
II.4 Experimental demonstration on quantum device
@@
Figure 10: Success rate of the QCNN during training on a quantum device. The experiment was conducted in 12 groups, each spaced hours apart to assess drift error. Each group involved $10^{4}$ measurement repeats to determine the success probability. Error bars indicated by the shaded region represent the standard error across different groups.
@@
Figure 11: Error mitigation ansatz. (a)&(b): Two different methods for boosting errors in the circuit: (a) adding single-qubit Pauli errors; (b) replacing $U$ with $U\left(U^{\dagger}U\right)^{r}$. (c) Error mitigation according to input fidelity. Linear regression suggests $P_{1}\approx 67\pm 1\%$ in the limit of a perfect input state. The red and blue dots correspond to Monte Carlo sampling of single-qubit Pauli errors (averaged over $10^{3}$ noise realizations) and repeating $U^{\dagger}U$, respectively. (d) Error mitigation according to error rate in the second approach. Linear regression using data prior to saturation suggests $P_{1}\approx 63\pm 1\%$ in the limit of a perfect input state.
II.5 Generalizations to other models
@@
Figure 12: Additional marked states in the PXP model. (a) Overlap between the eigenstates and the $|Z_{2}\rangle$ state, with red crosses marking the states identified by the QCNN. The yellow squares are the eigenstates of quasimodes in the symmetric subspace $\mathcal{K}$.(b) Weights of each eigenstates within the symmetric subspace $\mathcal{K}$. (c) Density of QCNN-identified states based in (a). (d) Density of energy differences between the QCNN-identified states in (a). In both (c) and (d), only states with non-zero overlap with the $|Z_{2}\rangle$ state are counted, with smoothing applied using Gaussian broadening of each points set to $20.1\Omega/$ 2.
@@
In this section, we extend our QCNN-based approach to two additional models that host QMBSs. We begin with the PXP model, utilizing training data that include both analytically solvable and numerically identified scar states. Specifically, we incorporate the four exact scar states reported in Ref. [32] and consider the states with the largest overlap with the $|Z_{2}\rangle$ across various energy windows as scar states [31]. Notably, the QCNN identifies many additional states as potentially non-thermal states, as shown in Fig. 12(a). Some of these states exhibit smaller overlaps with $|Z_{2}\rangle$, remaining hidden within the chaotic spectrum. Inspired by Ref. [43], a symmetric subspace $\mathcal{K}$ can be constructed, which exhibits regular motion. The quasimodes within this subspace can be viewed as approximations of certain eigenstates. The subspace $\mathcal{K}$ is spanned by the basis states
$|n_{1},n_{2}\rangle=\frac{1}{\sqrt{\mathcal{N}_{n_{1},n_{2}}}}\sum_{x\in(n_{1},n_{2})}|x\rangle\leavevmode\nobreak\ ,$ (10)
where $x$ represents the binary configuration of the spin chain, with the constraint that neighboring spins cannot both be in the "1" state. Here, $n_{1}$ and $n_{2}$ denote the total number of "1"s at odd and even positions, respectively, and $\mathcal{N}_{n_{1},n_{2}}$ is the normalization factor. The Hamiltonian in this subspace is expressed as $\langle n_{1},n_{2}|H_{\rm xorX}|n_{1}^{\prime},n_{2}^{\prime}\rangle$. The eigenstates of quasimodes in this subspace are shown as yellow squares in Fig. 12(a), and these quasimodes closely align with some of the marked states. In Fig. 12(b), we plot the probability distribution of eigenstates within the subspace $\mathcal{K}$. Notably, the marked states have significant components within this subspace, demonstrating that the QCNN effectively learns hidden properties of the quasimodes without prior knowledge. Additionally, the QCNN not only identifies the top band in Fig. 12(a), which is near the quasimodes, but also marks a second, lower band. This band has attracted considerable interest, though it still lacks a theoretical explanation.
The dynamics of these states are governed by their energy spectrum. The marked states exhibit energies similar to those of the $Z_{2}$ tower states. These states form distinct energy towers, as depicted in Fig. 12(c). Their energies are approximately equally spaced, as shown in Fig. 12(d), indicating that their linear superposition can lead to stable oscillations.
The perturbations in the PXP model exhibit distinct behaviors. In cases where perturbations enhance quantum many-body scars [44, 45, 46], we observe a reduction in the number of non-thermal states identified by the QCNN. Under a uniform magnetic field perturbation, the QCNN additionally marks the second-highest energy band, alongside the top band, which is generally recognized as hosting scar states. Results for various perturbations are presented in Appendix C.
@@
Figure 13: Additional marked states in the ar-coupling Ising SSH model. (a) Overlap between the eigenstates and the $Z_{1001}$ state. The red crosses are states marked by QCNN. (b) Density of QCNN-identified states based in (a). (c) Density of energy differences between the QCNN-identified states in (a). In both (b) and (c), only states with non-zero overlap with the $|Z_{1001}\rangle$ state are counted, with smoothing applied using Gaussian broadening of each points set to $0.25J_{\rm e}$.
@@
III DISCUSSIONS
Recent advances in the use of QML to investigate scar states have garnered considerable interest, resulting in a variety of related research. Here, we discuss differences between their studies and our approach. For instance, Ref. [28] introduces a general classifier using quantum variational autoencoders, where each eigenstate is assigned its own learning circuit. This approach can be computationally expensive, with costs scaling exponentially as Hilbert spaces grow in size. In contrast, our method employs a single circuit to learn the common properties of known scar states, enabling the discovery of other similar states. This makes our scheme more computationally feasible. Meanwhile, Ref. [47] employs classical machine learning to detect single-body scar in billiard systems, given the classical description of eigenstates, but this methodology does not extend to many-body scars. Ref. [29] uses classical machine learning and dimensionality reduction techniques to minimize the distance among scar states, but this can lead to loss of quantum information. Although recent work shows that certain QCNNs can be simulated efficiently on classical systems [48], our QCNN differs in two key ways: (1) Our QCNN is enhanced with additional universal layers in front of the original QCNN. We anticipate that using a small number of these layers (scaling logarithmically or linearly with system size) will make classical simulation hard. (2) Preparing eigenstates of many-body systems remains a computationally challenging task. Scaling up system sizes will still rely on quantum hardware, complemented by techniques such as the variational eigensolver [49] for efficient eigenstate preparation.
The convolutional layers in QCNNs are designed to detect local patterns and correlations in quantum states, making them particularly effective for identifying scar states, which exhibit sub-volume law entanglement entropy. Scar states are well-represented by matrix product states (MPS) [50, 32, 51], which can be efficiently generated by tree-structured quantum circuits [52]. Given that QCNNs have an inverse tree structure, they are inherently well-suited for capturing the characteristics of scar states. Furthermore, the pooling operation in QCNNs parallels the coarse-graining step in the renormalization group (RG), simplifying the system while retaining essential information. Both QCNNs and RG employ multi-scale analysis to extract key features, albeit in different contexts [27]. A deeper exploration of this correspondence and its underlying mechanisms remains an open avenue for further investigation.
@@ III DISCUSSIONS
IV METHODS
IV.1 Enhancement of Quantum Convolutional Neural Network
We further enhance the QCNN to improve its ability to classify quantum states with greater precision. Specifically, we introduce additional convolution layers after each existing convolution layer to enable the network to capture correlations over longer distances, analogous to the effect of using larger kernels in classical convolutional neural networks. The performance with different numbers of convolutional layers is shown in Fig. 14. To tackle more complex classification tasks, we introduce a few general layers of universal quantum gates to preprocess the input data. This adjustment helps handle finite-size effect and outliers which breaks translational-symmetry, such as disorders or boundaries. Furthermore, we incorporate an ancillary qubit initialized in the $|0\rangle$ state, analogous to the zero padding at the boundary pixels in classical convolutional neural networks. This allows us to perform classification operations without altering the size of the Hilbert space (see Appendix A).
IV.2 Training with Scar States
Figure 14: Training the xorX model with varying numbers of trainable gate parameters, denoted as $n_{\theta}$. The rate of marked states $r_{m}$ is the ratio of marked states to the total dimension of the Hilbert space. Each point represents a distinct set of circuit parameters and a corresponding training iteration. The orange dashed line indicates the rate of marked (scar) states in the input data. The system consists of 12 spins. The number of trainable parameters is given by $n_{\theta}=12+3(9n_{l}+3)$, where $n_{l}$ is the number of convolutional layers preceding the pooling layers.
$\mathcal{L}$ $=$ $\frac{1}{d}\sum_{i=0}^{d-1}\left|y_{i}-q_{i}\right|\leavevmode\nobreak\ ,$ (11)
where $d$ is the total number of input states, and $q_{i}$ denotes the probability of the output qubit being in |$\ket{1}$ state. This $q_{i}$ also corresponds to the single-shot measurement accuracy for identifying exact scar states. The enhanced QCNN is expected to distinguish scar states from thermal states after sufficient training with enough number of layers. After training, we evaluate the QCNN’s performance by testing it on the eigenstates of the Hamiltonian. For convenience, we consider an eigenstate to be “marked" if the output $q_{i}$ exceeds 50%, matching the rate of scar states in the training set. Experimentally, this classification is achieved through a majority vote across multiple trials and measurements [37]. In this context, achieving zero loss indicates perfect classification, although zero classification error does not necessarily imply the loss function fully decays to zero. The loss function also reflects the model’s robustness to noise.
Acknowledgements.
The authors acknowledge discussions with Zlatko Papic on quasi-modes in the PXP model. J.-J.F., B.Z. and Q.Z. acknowledge support from University of Southern California. Z.-C. Y. acknowledges support from Peking University. This research was conducted using IBM Quantum Systems provided through USC’s IBM Quantum Innovation Center. J.-J.F. and Q.Z. proposed the study of scar states using quantum machine learning. Q.Z. designed the framework of quantum machine learning. J.-J.F. performed numerical analyses and experiments, with support from B.Z., under the supervision of Q.Z. J.-J.F. derived the analytical theory. Z.-C.Y. provided insights on quantum many body scars. J.-J.F. and Q.Z. wrote the manuscript, with inputs and contributions from all authors.
@@ Appendix A Details on the QCNN
$U|A_{i}\rangle$ $=$ $|a_{i}\rangle\otimes|0\rangle\leavevmode\nobreak\ ,$ (12)
$U|B_{j}\rangle$ $=$ $|b_{j}\rangle\otimes|1\rangle\leavevmode\nobreak\ ,$ (13)
@@
$U|A_{i}\rangle\otimes|0\rangle$ $=$ $|\alpha_{i}\rangle\otimes|0\rangle\leavevmode\nobreak\ ,$ (14)
$U|B_{j}\rangle\otimes|0\rangle$ $=$ $|\beta_{j}\rangle\otimes|1\rangle\leavevmode\nobreak\ ,$ (15)
@@ Appendix B Details of experiments
Several methods exist for generating exact scar states [55, cenedese2024shallow]. For producing the first exact scar state, we employ a more convenient approach. The procedure for generating the first scar state is outlined in Fig. 17. All the qubits are initialized to $|0\rangle$. The last qubit $q_{n-1}$ is flipped to $|1\rangle$ using a not gate (X). An equal superposition is created on the ancilla qubit with a Hadamard gate (H). This is followed by a controlled-SWAP gate, which generates a superposition of spins at different positions $\left(|0_{\rm a}\rangle\otimes|100\cdots\rangle+|1_{\rm a}\rangle\otimes|\cdots 010\cdots\rangle\right)/\sqrt{2}$. Next, a CNOT gate is applied to restore the ancilla qubit, resulting in $|0_{\rm a}\rangle\otimes\left(|100\cdots\rangle+|\cdots 010\cdots\rangle\right)/\sqrt{2}$. Uniform distribution along the chain requires $O(n)$ equal divisions. Finally, a Z gate is applied to adjust the phase.
@@
${\rm Tr}\left\{|\mathscr{S}_{1}\rangle\langle\mathscr{S}_{1}|\rho_{\varepsilon}\right\}$ $=$ ${\rm Tr}\large\{U|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|U^{\dagger}U_{\varepsilon}|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|U_{\varepsilon}^{\dagger}\large\}\leavevmode\nobreak\ ,$
@@
$X_{n-1}U^{\dagger}U_{\varepsilon}X_{n-1}\approx U^{\dagger}U_{\varepsilon}\leavevmode\nobreak\ .$ (17)
@@
$\approx$ ${\rm Tr}\large\{|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|H_{\rm s}U_{\varepsilon}X_{n-1}|\mathscr{S}_{0}\rangle\langle\mathscr{S}_{0}|X_{n-1}U_{\varepsilon}^{\dagger}H_{\rm s}\large\}\leavevmode\nobreak\ .$
@@
Appendix C PXP Model Under Perturbations
Figure 18: (color online) (a) The eigenstates of $H_{\rm PXPZ}$ with perturbation strength $\lambda=0.05$. (b) The eigenstates of $H_{\rm sta}$ with perturbation strength $\lambda=0.35$. (c) The eigenstates of $H_{\rm Z}$ with perturbation strength $\lambda=0.7$.
The perturbative PXP models have garnered significant interest in the community due to their ability to alter both the spectral distribution and quench dynamics. In this section, we numerically evaluate the performance of the QCNN with continua boundary condition under Hamiltonian perturbations. Specifically, we consider three distinct types of perturbations.
Scars can be enhanced by incorporating higher-order corrections into the Hamiltonian, which in turn strengthens the revival. This modification can be expressed as [44, 45]
$H_{\rm PXPZ}$ $=H_{\rm PXP}$
$-\lambda\sum_{i=0}^{n}\left(\sigma_{i-2}^{z}P_{i-1}^{0}\sigma_{i}^{x}P_{i+1}^{0}+P_{i-1}^{0}\sigma_{i}^{x}P_{i+1}^{0}\sigma_{i+2}^{z}\right)\leavevmode\nobreak\ ,$
where $\lambda$ is the perturbation strength.
A staggered magnetic field induces the confinement of excitations, effectively isolating the antiferromagnetic-like states from other states. This separation also enhances the revival strength. The corresponding Hamiltonian is [46]
$H_{\rm sta}=H_{\rm PXP}+\lambda\sum_{i=0}^{n}(-1)^{i}\sigma_{i}^{z}\leavevmode\nobreak\ .$ (20)
In the presence of a chemical potential, the reference state shifts to the ferromagnetic state $|F\rangle=|0\rangle^{\otimes n}$. The corresponding Hamiltonian is [57, 58]
$H_{\rm Z}=H_{\rm PXP}+\lambda\sum_{i=0}^{n}\sigma_{i}^{z}\leavevmode\nobreak\ .$ (21)
We train the QCNN using the scar states in the top band for each perturbative case. In all cases, the QCNN identifies most of the scar states in the top band, as shown in Fig. 18. For the revival-enhanced cases in Fig. 18 (a) and (b), fewer additional states are detected, suggesting that fewer scar-like states exist in these perturbative systems. This indicates that the dispersion is suppressed. For the reference state switch case shown in Fig. 18 (c), the QCNN also identifies states in the second band below the top band, implying that these states share similar properties with the scar states. A more detailed and in-depth investigation is needed in future studies.